OpenStudy (anonymous):

Our professor never showed us how to work this out. I checked the internet and was told there was no solution, but how would you work this out? 5(x-1)+3squareroot(x-1)-2=0

OpenStudy (ybarrap):

5(x - 1) - 3(x - 1)^(1/2) - 2 = 0 so x = 2.

OpenStudy (ybarrap):

5 ( 2 - 1) - 3 (2 - 1)^(1/2) - 2 = 5 (1) - 3 (1)^(1/2) -2 = 5 - 3 - 2 = 0

OpenStudy (ybarrap):

Solved by inspection

OpenStudy (anonymous):

Okay, thank you so much!